A rectangle has a length of 12 m and a width of 400 cm. What is the perimeter of the rectangle?
Answer by jim_thompson5910(35256)

400 cm = 4 m, so the width is 4 meters

-------------------------------------------------------


P = 2*(L + W)


P = 2*(12 + 4)


P = 2*16


P = 32


The perimeter is 32 meters (or 32*100 = 3200 cm)